What must be the ratio of F- concentration to HF to make a buffer with a pH of 4? pKa = 3.2

a. 0.16

b. 2.4

c. 6.3

d.8.1


1
Expert's answer
2022-01-19T11:32:01-0500

pH = 4

[H] =10-pH

= 10-4 = 1.0x10-4

Also;

[H] "=\dfrac{nHF x Ka}{nF-}" Since pKa = -LogKa

Ka = 10-pKa = 10-3.2


1.0x10-4 "=" "\dfrac{nHF x 10^-3.2}{nF^-}"


"\dfrac{nHF}{nF^-}" = 0.158 = 0.16